OpenStudy (anonymous):

A kilobyte is 2^10 bytes. A gigabyte is 2^30 bytes. Use exponent rules to find how many kilobytes are in a gigabyte. Give your answer in exponent form.

OpenStudy (anonymous):

when dividing powers, subtract exponents. there are 2^30/2^10 = 2^(30 - 10) = 2^20 kB per gB.

OpenStudy (anonymous):

So 2^20 kilobytes per gb?

OpenStudy (anonymous):

1 gigabyte = 2^30 bytes = 2^20 * 2^10 bytes = 2^20kilobytes* 1 kilobytes = 2^20 kb

OpenStudy (anonymous):

\[2^{20}\] gb

OpenStudy (anonymous):

Thanks guys. :)